OpenStudy (anonymous):

Regan is trying to find the equation of a quadratic that has a focus of (−2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (anonymous):

(-2,5) @surry99 can you help me with this please

OpenStudy (anonymous):

but its negative

OpenStudy (surry99):

Are you referring to the -2?

OpenStudy (anonymous):

yes i am

OpenStudy (surry99):

no problem...if you look at the general formula... (x-a)^2 + b^2 - c^2 = 2(b-c)y in your problem the focus is (-2, 5) which is (a, b) so..a = -2 and b =5 the directrix y =13 which is y=c...so c =13. So plug these values in to the general formula and simplify. (x - -2)^2 + 5^2 - 13^2 = 2(5-13)y... and (x--2) = (x+2) so... (x+2)^2 +5^2 -13^2 = 2(5-13)y...now expand and simplify

OpenStudy (anonymous):

thanks